# Divergences induced by dual subtractive and divisive normalizations of exponential families and their convex deformations

### 20 Dec 2023 v1

@@ 4 Divergences related to the partition function
The Hellinger divergence [3] between two positive potentially unnormalized densities ${\tilde{p}}$ and ${\tilde{q}}$ is defined by
$D_{H}({\tilde{p}},{\tilde{q}})$ $=$ $\frac{1}{2}\int(\sqrt{{\tilde{p}}}-\sqrt{{\tilde{q}}})^{2}\,\mathrm{d}\mu,$
@@
Notice that the Hellinger divergence can be interpreted as the integral of the difference between the arithmetic mean $A({\tilde{p}},{\tilde{q}})=\frac{{\tilde{p}}+{\tilde{q}}}{2}$ minus the geometric mean $G({\tilde{p}},{\tilde{q}})=\sqrt{{\tilde{p}}{\tilde{q}}}$ of the densities: $D_{H}({\tilde{p}},{\tilde{q}})=\int(A({\tilde{p}},{\tilde{q}})-G({\tilde{p}},{\tilde{q}}))\mathrm{d}\mu$ This also proves that $D_{H}({\tilde{p}},{\tilde{q}})\geq 0$ since $A\geq G$.
@@
$D_{H}({\tilde{p}}_{\theta_{1}},{\tilde{p}}_{\theta_{2}})=\frac{Z(\theta_{1})+Z(\theta_{2})}{2}-Z\left(\frac{\theta_{1}+\theta_{2}}{2}\right)=J_{Z}(\theta_{1},\theta_{2}),$ (12)
@@
$D_{\alpha}({\tilde{p}}:{\tilde{q}})=\left\{\begin{array}[]{ll}\frac{1}{\alpha(1-\alpha)}\int\left(\alpha{\tilde{p}}+(1-\alpha){\tilde{q}}-{\tilde{p}}^{\alpha}{\tilde{q}}^{1-\alpha}\right)\mathrm{d}\mu,&\alpha\not\in\{0,1\}\\ D_{\mathrm{KL}}^{*}({\tilde{p}}:{\tilde{q}})=D_{\mathrm{KL}}({\tilde{q}}:{\tilde{p}})&\alpha=0,\\ 4\,D_{H}({\tilde{p}},{\tilde{q}})&\alpha=\frac{1}{2},\\ D_{\mathrm{KL}}({\tilde{p}}:{\tilde{q}})&\alpha=1.\end{array}\right.$
We have $D_{\alpha}^{*}({\tilde{p}}:{\tilde{q}})=D_{\alpha}({\tilde{q}}:{\tilde{p}})=D_{1-\alpha}({\tilde{p}}:{\tilde{q}})$.
@@
$H^{\times}({\tilde{p}}:{\tilde{q}})$ $=$ $Z_{p}\,H^{\times}(p:q)-Z_{p}\log Z_{q}+Z_{q}-1,$
$H({\tilde{p}})$ $=$ $-Z_{p}\,H(p)-Z_{p}\log Z_{p}+Z_{p}-1,$
@@
$B_{Z}(\theta_{2}:\theta_{1})=Z(\theta_{1})\,B_{F}(\theta_{2}:\theta_{1})+Z(\theta_{1})\log\frac{Z(\theta_{1})}{Z(\theta_{2})}+Z(\theta_{2})-Z(\theta_{1}).$ (19)
Let $D_{\mathrm{skl}}(a:b)=a\log\frac{a}{b}+b-a$ be the scalar KLD. Then we have: $B_{Z}(\theta_{2}:\theta_{1})=Z(\theta_{1})\,B_{F}(\theta_{2}:\theta_{1})+D_{\mathrm{skl}}(Z(\theta_{1}):Z(\theta_{2}))$.
Let us illustrate Proposition 5 with an example:
Example 1
@@
$D_{H}({\tilde{p}}_{\sigma^{2}_{1}}:{\tilde{p}}_{\sigma^{2}_{2}})=J_{Z}(\theta_{1},\theta_{2})=\sqrt{\frac{\pi}{2}}(\sigma_{1}+\sigma_{2})-2\sqrt{\pi}\frac{\sigma_{1}\sigma_{2}}{\sqrt{\sigma_{1}^{2}+\sigma_{2}^{2}}},$
@@
5 Deforming convex functions and dually flat spaces
@@ 5.1 Comparative convexity
The log-convexity can be interpreted as a special case of comparative convexity with respect to a pair $(M,N)$ of comparable weighted means [17]: A function $Z$ is $(M,N)$-convex iff for $\alpha\in[0,1]$, we have
$Z(M(x,y;\alpha,1-\alpha))\leq N(Z(x),Z(y);\alpha,1-\alpha),$
and strictly $(M,N)$-convex iff we have strict inequality for $\alpha\in(0,1)$ and $x\not=y$. Log-convexity corresponds to $(A,G)$-convexity, i.e., convexity with respect to the weighted arithmetic and geometric means: $A(x,y;\alpha,1-\alpha)=\alpha x+(1-\alpha)y$ and $G(x,y;\alpha,1-\alpha)=x^{\alpha}y^{1-\alpha}$.
A quasi-arithmetic mean is defined for a continuous and strictly increasing function $h$ by
@@
Proposition 6 ([1, 25])
A function $Z(\theta)$ is strictly $(M_{\rho},M_{\tau})$-convex iff the function $\tau\circ Z\circ\rho^{-1}$ is strictly convex.
Since log-convexity is $(A=M_{\mathrm{id}},G=M_{\log})$-convexity, a function $Z$ is convex iff $\log\circ Z\circ\mathrm{id}^{-1}=\log\circ Z$ is strictly convex. We have
@@
Starting from a convex function $F$, we can deform the function to $Z$ using strictly monotone functions $\tau$ and $\rho$. For a $(M_{\rho},M_{\tau})$-convex function $Z(\theta)$ which is also strictly convex, we can define a pair of Bregman divergences $B_{Z}$ and $B_{F}$ with $F(\theta)=\tau(Z(\rho^{-1}(\theta)))$ and a pair of skewed Jensen divergences.
Example 2
Consider $\rho=h_{p_{1}}$, $\tau=h_{p_{2}}$ for $p_{1},p_{2}\in\mathbb{R}$, and the convex partition function $Z(\theta)$ of an exponential family. Then the function $F(\theta)=\tau(Z(\rho^{-1}(\theta)))$ has second derivative:
$F^{\prime\prime}(\theta)=Z\left(\left(p_{1}\,x+1\right)^{{{1}\over{p_{1}}}}\right)^{p_{2}-1}\,\left({{d^{2}}\over{d\,x^{2}}}\,Z\left(\left(p_{1}\,x+1\right)^{{{1}\over{p_{1}}}}\right)\right)+\left(p_{2}-1\right)\,Z\left(\left(p_{1}\,x+1\right)^{{{1}\over{p_{1}}}}\right)^{p_{2}-2}\,\left({{d}\over{d\,x}}\,Z\left(\left(p_{1}\,x+1\right)^{{{1}\over{p_{1}}}}\right)\right)^{2}.$
Since $Z$ is strictly convex and positive, we have $F$ that is strictly convex when $p_{2}\geq 1$ for any value of $p_{1}\in\mathbb{R}$. Thus we can build a $2$-parameter deformation of the partition function which remains convex:
$Z_{p_{1},p_{2}}=h_{p_{2}}\circ Z\circ h_{p_{1}}^{-1},\quad p_{1}\geq 1,p_{2}\in\mathbb{R}.$
This yields corresponding families of Bregman divergences which can be interpreted as statistical divergences, and skew Jensen divergences.

### 18 Jan 2024 v2

@@ 4 Divergences related to the partition function
The squared Hellinger distance [4] between two positive potentially unnormalized densities ${\tilde{p}}$ and ${\tilde{q}}$ is defined by
$D_{H}^{2}({\tilde{p}},{\tilde{q}})$ $=$ $\frac{1}{2}\int(\sqrt{{\tilde{p}}}-\sqrt{{\tilde{q}}})^{2}\,\mathrm{d}\mu,$
@@
Notice that the Hellinger divergence can be interpreted as the integral of the difference between the arithmetic mean $A({\tilde{p}},{\tilde{q}})=\frac{{\tilde{p}}+{\tilde{q}}}{2}$ minus the geometric mean $G({\tilde{p}},{\tilde{q}})=\sqrt{{\tilde{p}}{\tilde{q}}}$ of the densities: $D_{H}^{2}({\tilde{p}},{\tilde{q}})=\int(A({\tilde{p}},{\tilde{q}})-G({\tilde{p}},{\tilde{q}}))\mathrm{d}\mu$ This also proves that $D_{H}({\tilde{p}},{\tilde{q}})\geq 0$ since $A\geq G$. The Hellinger distance $D_{H}$ satisfies the metric axioms of distances.
@@
$D_{H}^{2}({\tilde{p}}_{\theta_{1}},{\tilde{p}}_{\theta_{2}})=\frac{Z(\theta_{1})+Z(\theta_{2})}{2}-Z\left(\frac{\theta_{1}+\theta_{2}}{2}\right)=J_{Z}(\theta_{1},\theta_{2}),$ (14)
@@
$D_{\alpha}({\tilde{p}}:{\tilde{q}})=\left\{\begin{array}[]{ll}\frac{1}{\alpha(1-\alpha)}\int\left(\alpha{\tilde{p}}+(1-\alpha){\tilde{q}}-{\tilde{p}}^{\alpha}{\tilde{q}}^{1-\alpha}\right)\mathrm{d}\mu,&\alpha\not\in\{0,1\}\\ D_{\mathrm{KL}}^{*}({\tilde{p}}:{\tilde{q}})=D_{\mathrm{KL}}({\tilde{q}}:{\tilde{p}})&\alpha=0,\\ 4\,D_{H}^{2}({\tilde{p}},{\tilde{q}})&\alpha=\frac{1}{2},\\ D_{\mathrm{KL}}({\tilde{p}}:{\tilde{q}})&\alpha=1.\end{array}\right.$
We have $D_{\alpha}^{*}({\tilde{p}}:{\tilde{q}})=D_{\alpha}({\tilde{q}}:{\tilde{p}})=D_{1-\alpha}({\tilde{p}}:{\tilde{q}})$. and the $\alpha$-divergences are homogeneous divergences of degree $1$: For all $\lambda>0$, $D_{\alpha}(\lambda{\tilde{q}}:\lambda{\tilde{p}})=\lambda\,D_{\alpha}({\tilde{q}}:{\tilde{p}})$. Moreoever, since $\alpha{\tilde{p}}+(1-\alpha){\tilde{q}}-{\tilde{p}}^{\alpha}{\tilde{q}}^{1-\alpha}$ can be expressed as the difference of the weighted arithmetic minus the geometric means $A({\tilde{p}},{\tilde{q}};\alpha;1-\alpha)-G({\tilde{p}},{\tilde{q}};\alpha;1-\alpha)$, it follows from the arithmetic-geometric mean inequality that we have $D_{\alpha}({\tilde{p}}:{\tilde{q}})\geq 0$.
@@
Proof.
For $\alpha\not\in\{0,1\}$, consider
$D_{\alpha}({\tilde{p}}_{\theta_{1}}:{\tilde{p}}_{\theta_{2}})=\frac{1}{\alpha(1-\alpha)}\int\left(\alpha{\tilde{p}}_{\theta_{1}}+(1-\alpha){\tilde{p}}_{\theta_{2}}-{\tilde{p}}_{\theta_{1}}^{\alpha}{\tilde{p}}_{\theta_{2}}^{1-\alpha}\right)\mathrm{d}\mu.$
We have $\int\alpha{\tilde{p}}_{\theta_{1}}\,\mathrm{d}\mu=\alpha Z(\theta_{1})$, $\int(1-\alpha){\tilde{p}}_{\theta_{2}}\,\mathrm{d}\mu=(1-\alpha)Z(\theta_{2})$, and $\int{\tilde{p}}_{\theta_{1}}^{\alpha}{\tilde{p}}_{\theta_{2}}^{1-\alpha}\mathrm{d}\mu=\int{\tilde{p}}_{\alpha\theta_{1}+(1-\alpha)\theta_{2}}\mathrm{d}\mu=Z(\alpha\theta_{1}+(1-\alpha)\theta_{2})$. It follows that $D_{\alpha}({\tilde{p}}_{\theta_{1}}:{\tilde{p}}_{\theta_{2}})=\frac{1}{\alpha(1-\alpha)}J_{Z,\alpha}(\theta_{1}:\theta_{2})==J_{Z,\alpha}^{s}(\theta_{1}:\theta_{2})$. ∎
@@
$H^{\times}({\tilde{p}}:{\tilde{q}})$ $=$ $Z_{p}\,H^{\times}(p:q)-Z_{p}\log Z_{q}+Z_{q}-1,$ (21)
$H({\tilde{p}})$ $=$ $Z_{p}\,H(p)-Z_{p}\log Z_{p}+Z_{p}-1,$ (22)
@@
$B_{Z}(\theta_{2}:\theta_{1})$ $=$ $Z(\theta_{1})\,B_{F}(\theta_{2}:\theta_{1})+Z(\theta_{1})\log\frac{Z(\theta_{1})}{Z(\theta_{2})}+Z(\theta_{2})-Z(\theta_{1}),$ (23)
$=$ $\exp(F(\theta_{1}))\,B_{F}(\theta_{2}:\theta_{1})+(\exp F(\theta_{1}))(F(\theta_{1})-F(\theta_{2}))+\exp(F(\theta_{2}))-\exp(F(\theta_{1})).$ (24)
Let $D_{\mathrm{skl}}(a:b)=a\log\frac{a}{b}+b-a$ be the scalar KLD for $a>0$ and $b>0$. Then we rewrite Eq. 20 as
$D_{\mathrm{KL}}({\tilde{p}}:{\tilde{q}})=Z_{p}\,D_{\mathrm{KL}}(p:q)+D_{\mathrm{skl}}(Z_{p}:Z_{q}),$
and we have:
$B_{Z}(\theta_{2}:\theta_{1})=Z(\theta_{1})\,B_{F}(\theta_{2}:\theta_{1})+D_{\mathrm{skl}}(Z(\theta_{1}):Z(\theta_{2})).$
The KLD between unnormalized densities ${\tilde{p}}$ and ${\tilde{q}}$ with support $\mathcal{X}$ can also be written as a definite integral of a scalar Bregman divergence:
$D_{\mathrm{KL}}({\tilde{p}}:{\tilde{q}})=\int_{\mathcal{X}}D_{\mathrm{skl}}({\tilde{p}}(x):{\tilde{q}}(x))\,\mathrm{d}\mu(x)=\int_{\mathcal{X}}B_{f_{\mathrm{skl}}}({\tilde{p}}(x):{\tilde{q}}(x))\,\mathrm{d}\mu(x),$
where $f_{\mathrm{skl}}(x)=x\log x-x$. Since $B_{f_{\mathrm{skl}}}(a:b)\geq 0\forall a>0,b>0$, we deduce that $D_{\mathrm{KL}}({\tilde{p}}:{\tilde{q}})\geq 0$ with equality iff ${\tilde{p}}(x)={\tilde{q}}(x)$ $\mu$-almost everywhere.
Notice that $B_{Z}(\theta_{2}:\theta_{1})=Z(\theta_{1})\,B_{F}(\theta_{2}:\theta_{1})+D_{\mathrm{skl}}(Z(\theta_{1}):Z(\theta_{2}))$ can be interpreted as the sum of two divergences: a conformal Bregman divergence with a scalar Bregman divergence.
Remark 5.
Consider the KLD between the normalized $p_{\theta_{1}}$ and unnormalized ${\tilde{p}}_{\theta_{2}}$ densities of a same exponential family. We have
$D_{\mathrm{KL}}(p_{\theta_{1}}:{\tilde{p}}_{\theta_{2}})$ $=$ $B_{F}(\theta_{2}:\theta_{1})-\log Z(\theta_{2})+Z(\theta_{2})-1,$
$=$ $Z(\theta_{2})-1-F(\theta_{1})-{\left\langle\theta_{2}-\theta_{1},\nabla F(\theta_{2})\right\rangle},$
$=$ $B_{Z-1,F}(\theta_{2}-\theta_{1}).$ (26)
Divergence $B_{Z-1,F}$ is a duo Bregman pseudo-divergence [31]:
$B_{F_{1},F_{2}}(\theta_{1}:\theta_{2})=F_{1}(\theta_{1})-F_{2}(\theta_{2})-{\left\langle\theta_{1}-\theta_{2},\nabla F_{2}(\theta_{2})\right\rangle},$
for $F_{1}$ and $F_{2}$ two strictly convex and smooth functions such that $F_{1}\geq F_{2}$. Indeed, we check that both generators $F_{1}(\theta)=Z(\theta)-1$ and $F_{2}(\theta)=F(\theta)$ are Bregman generators, and we have $F_{1}(\theta)\geq F_{2}(\theta)$ since $e^{x}\geq x+1$ for all $x$ (with equality when $x=0$), i.e., $Z(\theta)-1\geq F(\theta)$.
More generally, the $\alpha$-divergences between $p_{\theta_{1}}$ and ${\tilde{p}}_{\theta_{2}}$ can be written as
$D_{\alpha}(p_{\theta_{1}}:{\tilde{p}}_{\theta_{2}})=\frac{1}{\alpha(1-\alpha)}\left(\alpha Z(\theta_{1})+(1-\alpha)-\frac{Z(\alpha\theta_{1}+(1-\alpha)\theta_{2})}{Z(\theta_{2})}.\right),$ (27)
and the (signed) $\alpha$-skewed Bhattacharyya distances are given by
$D_{B,\alpha}(p_{\theta_{1}}:{\tilde{p}}_{\theta_{2}})=\log Z(\theta_{2})-\log Z(\alpha\theta_{1}+(1-\alpha)\theta_{2}).$
Let us illustrate Proposition 5 with some examples:
Example 1.
Consider the family of exponential distributions =$\mathcal{E}=\{p_{\lambda}(x)=1_{x\geq 0}\,\lambda\exp(-\lambda x)\}$. $\mathcal{E}$ is an exponential family with natural parameter $\theta=\lambda$ and parameter space $\Theta=\mathbb{R}{>0}$, and sufficient statistic $t(x)=-x$. The partition function is $Z(\theta)=\frac{1}{\theta}$ with $Z^{\prime}(\theta)=-\frac{1}{\theta^{2}}$ and $Z^{\prime\prime}(\theta)=\frac{2}{\lambda^{3}}>0$. The cumulant function is $F(\theta)=\log Z(\theta)=-\log\theta$. with moment parameter $\eta=E_{p_{\lambda}}[t(x)]=F^{\prime}(\theta)=-\frac{1}{\theta}$. The $\alpha$-divergences between two unnormalized exponential distributions are:
$D_{\alpha}({\tilde{p}}_{\lambda_{1}}:{\tilde{p}}_{\lambda_{2}})=\left\{\begin{array}[]{ll}\frac{1}{\alpha(1-\alpha)}J_{Z,\alpha}(\theta_{1}:\theta_{2})=\frac{(\lambda_{1}-\lambda_{2})^{2})}{\alpha\lambda_{1}^{2}\lambda_{2}+(1-\alpha)\lambda_{1}\lambda_{2}^{2}}&\alpha\not\in\{0,1\}\\ D_{\mathrm{KL}}({\tilde{p}}_{\lambda_{2}}:{\tilde{p}}_{\lambda_{1}})=B_{Z}(\theta_{1}:\theta_{2})=\frac{(\lambda_{1}-\lambda_{2})^{2}}{\lambda_{1}\lambda_{2}^{2}}&\alpha=0,\\ 4\,J_{Z}(\theta_{1},\theta_{2})=\frac{(\lambda_{1}-\lambda_{2})^{2}}{2(\lambda_{1}\lambda_{2}^{2}+\lambda_{1}^{2}\lambda_{2})}&\alpha=\frac{1}{2},\\ D_{\mathrm{KL}}({\tilde{p}}_{\lambda_{1}}:{\tilde{p}}_{\lambda_{2}})=B_{Z}(\theta_{2}:\theta_{1})=\frac{(\lambda_{1}-\lambda_{2})^{2}}{\lambda_{2}\lambda_{1}^{2}}&\alpha=1\end{array}\right.$ (28)
Example 2.
@@
$D_{H}^{2}({\tilde{p}}_{\sigma^{2}_{1}}:{\tilde{p}}_{\sigma^{2}_{2}})=J_{Z}(\theta_{1},\theta_{2})=\sqrt{\frac{\pi}{2}}(\sigma_{1}+\sigma_{2})-2\sqrt{\pi}\frac{\sigma_{1}\sigma_{2}}{\sqrt{\sigma_{1}^{2}+\sigma_{2}^{2}}},$
@@
Consider now the family of univariate normal distributions
=$\mathcal{E}=\left\{p_{\mu,\sigma^{2}}(x)=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{1}{2}\left(\frac{x-\mu}{\sigma}\right)^{2}\right)\right\}.$
Let $\theta=\left(\theta_{1}=\frac{1}{\sigma_{2}},\theta_{2}=\frac{\mu}{\sigma^{2}}\right)$ and
$Z(\theta_{1},\theta_{2})=\sqrt{\frac{2\pi}{\theta_{1}}}\,\exp\left(\frac{1}{2}\frac{\theta_{2}^{2}}{\theta_{1}}\right).$
The unnormalized densities are ${\tilde{p}}_{\theta}(x)=\exp\left(-\frac{\theta_{1}x^{2}}{2}+x\theta_{2}\right)$. We have
$\nabla Z(\theta)=\left[\begin{array}[]{ll}\sqrt{\frac{\pi}{2}}\frac{(\theta_{1}+\theta_{2}^{2})\exp\left(\frac{\theta_{2}^{2}}{2\theta_{1}}\right)}{\theta_{1}^{\frac{5}{2}}}\\ \sqrt{2\pi}\frac{\theta_{2}\exp\left(\frac{\theta_{2}^{2}}{2\theta_{1}}\right)}{\theta_{1}^{\frac{3}{2}}}\end{array}\right].$
It follows that $D_{\mathrm{KL}}[{\tilde{p}}_{\theta}:{\tilde{p}}_{\theta^{\prime}}]=B_{Z}(\theta^{\prime}:\theta)$.
5 Deforming convex functions and their induced dually flat spaces
@@ 5.1 Comparative convexity
The log-convexity can be interpreted as a special case of comparative convexity with respect to a pair $(M,N)$ of comparable weighted means [29]:
A function $Z$ is $(M,N)$-convex if and only if for $\alpha\in[0,1]$, we have
$Z(M(x,y;\alpha,1-\alpha))\leq N(Z(x),Z(y);\alpha,1-\alpha),$ (29)
and strictly $(M,N)$-convex iff we have strict inequality for $\alpha\in(0,1)$ and $x\not=y$. Furthermore, a function $Z$ is (strictly) $(M,N)$-concave if $-Z$ is (strictly) $(M,N)$-convex.
Log-convexity corresponds to $(A,G)$-convexity, i.e., convexity with respect to the weighted arithmetic and geometric means defined respectively by $A(x,y;\alpha,1-\alpha)=\alpha x+(1-\alpha)y$ and $G(x,y;\alpha,1-\alpha)=x^{\alpha}y^{1-\alpha}$. Ordinary convexity is $(A,A)$-convexity.
A weighted quasi-arithmetic mean [26] (also called Kolmogorov-Nagumo mean [27]) is defined for a continuous and strictly increasing function $h$ by
@@
Proposition 6 ([1, 34]).
A function $Z(\theta)$ is strictly $(M_{\rho},M_{\tau})$-convex with respect to two strictly increasing smooth functions $\rho$ and $\tau$ if and only if the function $F=\tau\circ Z\circ\rho^{-1}$ is strictly convex.
Notice that the set of strictly increasing smooth functions form a non-abelian group with group operation the function composition, neutral element the identity function, and inverse element the functional inverse function.
Since log-convexity is $(A=M_{\mathrm{id}},G=M_{\log})$-convexity, a function $Z$ is strictly log-convex iff $\log\circ Z\circ\mathrm{id}^{-1}=\log\circ Z$ is strictly convex. We have
@@
Starting from a given convex function $F(\theta)$, we can deform the function $F(\theta)$ to obtain a function $Z(\theta)$ using two strictly monotone functions $\tau$ and $\rho$: $Z(\theta)=\tau^{-1}(F(\rho(\theta)))$.
For a $(M_{\rho},M_{\tau})$-convex function $Z(\theta)$ which is also strictly convex, we can define a pair of Bregman divergences $B_{Z}$ and $B_{F}$ with $F(\theta)=\tau(Z(\rho^{-1}(\theta)))$ and a corresponding pair of skewed Jensen divergences.
Thus we have the following generic deformation scheme:
[…]
In particular, when function $Z$ is deformed by strictly increasing power functions $h_{p_{1}}$ and $h_{p_{2}}$ for $p_{1}$ and $p_{2}$ in $\mathbb{R}$ as
$Z_{p_{1},p_{2}}=h_{p_{2}}\circ Z\circ h_{p_{1}}^{-1},$
$Z_{p_{1},p_{2}}$ is strictly convex when it is strictly $(M_{p_{1}},M_{p_{2}})$-convex, and thus induces corresponding Bregman and Jensen divergences.
Example 3.
Consider the partition function $Z(\theta)=\frac{1}{\theta}$ of the exponential distribution family ($\theta>0$ with $\Theta=\mathbb{R}_{>0}$). Let $Z_{p}(\theta)=(h_{p}\circ Z)(\theta)=\frac{\theta^{-p}-1}{p}$. Then we have $Z_{p}^{\prime\prime}(\theta)=(1+p)\frac{1}{\theta^{2+p}}>0$ when $p>-1$. Thus we can deform $Z$ smoothly by $Z_{p}$ while preserving convexity by ranging $p$ from $-1$ to +$\infty$+ ∞. We thus get a corresponding family of Bregman and Jensen divergences.
The proposed convex deformation by using quasi-arithmetic mean generators differs from the interpolation of convex functions using the technique of proximal average [8].